Cummins India Ltd has the following capital structure, which it considers optimal:

Debt 25%

Preference Shares 10%

Equity shares 65%

Total 100%

Applicable tax rate for the company is 25%. Risk free rate of return is 6%, average equity

market investment has expected rate of return of 12%. The company’s beta is 1.10.

Following terms would apply to new securities being issued as follows:

1. New preference can be issued at a face value of Rs. 100 per share, dividend and cost of

issuance will be Rs. 10 per share and Rs. 2 per share respectively.

2. Debt will bear an interest rate of 9%.

Calculate

a. component cost of debt, preference shares and equity shares assuming that the company

does not issue any additional equity shares. (5 Marks)

b. WACC. (5 Marks)



1
Expert's answer
2021-06-01T11:54:01-0400

a) Cost of Debt is given to be 10%


By issuing Preference shares, CP India Ltd can raise a net amount of Rs.100 -Rs.4 = Rs.96 and pay a dividend of Rs.8 annually


So, cost of preference shares to the company = Rs,8/Rs,96 = 0.0833 or 8.33%


Cost of equity (from CAPM) =risk free rate + beta × (Average equity return - risk free rate)


= 6%+1.10×(15%-6%)


=18.50%


b) WACC = weight of debt× cost of debt× (1-tax rate) + Weight of preference shares × cost of preference shares + weight of equity × cost of equity


=0.25×10%×(10.25)+0.15×8.33%+0.60×18.50%


=0.1443 or 14.43%